OpenStudy (anonymous):

State the domain and range of the following: 4x^2 +9y^2 = 36

OpenStudy (amistre64):

its eccliptic i beleive

OpenStudy (amistre64):

elliptic even

OpenStudy (amistre64):

one of the conics at any rate

OpenStudy (amistre64):

\[4x^2 +9y^2 = 36\] \[\frac{x^2}{36/4} +\frac{y^2}{36/9} = \frac{36}{36}\] \[(\frac{x}{3})^2 +(\frac{y}{2})^2 = 1\] the length of our x axis is 3 from center; and the y is 2 from center our center being (0,0)

OpenStudy (anonymous):

ecclictic as well.

OpenStudy (amistre64):

eclectic :)

OpenStudy (anonymous):

so it's a circle? I don't know what eclectic means

OpenStudy (amistre64):

its an oval; the rest of it can be defined on google if need be

OpenStudy (anonymous):

oh ok. I was reading it wrong. Thank you